OpenStudy (anonymous):

Solve. 5x2 - 3x = (5x2 - 7x) + (8x - 2) a: x = -2 b: x = 2 c: x = -1/2 d: x = 1/2

OpenStudy (mrnood):

the answer you chose for your LAST post was not correct

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can i just get the answer

OpenStudy (marihelenh):

\[5x^2-3x=5x^2-7x+8x-2\] Add like terms. \[5x^2-3x=5x^2+x-2\] Subtract 5x2 from both sides. \[-3x=x-2\] Subtract x from both sides. \[-4x=-2\] Divide both sides by -4. \[x=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

thanks